# Asteria: Semantic-Aware Cross-Region Caching for Agentic LLM Tool Access

### v1

@@ 5 Implementation and Discussion
Asteria is implemented in Python atop vLLM [63], which we use as the high-throughput serving layer to transparently intercept external tool calls and plug into existing agent applications (e.g., Search-R1 [32]). For co-location, we rely on CUDA MPS to let multiple processes share a single GPU context—avoiding OS-level context-switching overhead—and partition resources by setting CUDA_MPS_ACTIVE_THREAD_PERCENTAGE.
Judger accuracy. Our framework bridges the gap between caching and semantic knowledge reuse by providing a unified cache abstraction for LLM agents, and it leverages powerful LLM-based techniques from prior research to ensure strong accuracy. The semantic judger can be easily fine-tuned or replaced for specific workloads, so its accuracy can be improved with minimal effort when needed. Given the rapid progress of small LLMs, we find this sufficient for practical use and view the judger as a pluggable component. In practice, Asteria maintains high precision via strong semantic models and periodic recalibration, with misses falling back to live fetches to preserve correctness.
@@ 6.1 Experimental Setup
Testbed. We emulate a realistic cross-region deployment. The LLM agent and all Asteria components run on an on-premise H100 cluster; remote data services run in a separate region. For search agent, we use the public Google Cloud Search API whose per-request average latency ranges between 300–500ms depending on response length, consistent with prior work [40]. For coding, we use a self-deployed FAISS [16]-based RAG service [16] with an average 300 ms round trip.
Models and workloads. We evaluate two agentic scenarios: AI-powered search and code generation. The search agent is Search-R1-7B [33, 34] (post-trained from Qwen-2.5 7B [2]); the coding agent is Qwen-3 8B [3]. Asteria uses Qwen-3 0.6B models for embedding and semantic judging [60, 61]. Asteria’s internal embedding and semantic judging functions are powered by the lightweight 0.6B models from the Qwen-3 family [60, 61].
Search workloads. Following the Google Trend access patterns in subsection 2.3, we construct: (i) Skewed workload: extracting the Google-like head–tail access pattern in subsection 2.3, we instantiate four search benchmarks–Zilliz-GPT [68], HotpotQA [7], Musique [8], and 2Wiki-Multi-Hop [6]. For each dataset, we apply -means to the questions and keep 10 representative clusters, sampling 250 questions and constructing the skewed popularity per dataset (1000 total) and (ii) Trend-driven workload. We capture 12-hour Google Trends time series for four topics, map them to HotpotQA questions, and compress them into a 10-minute trace to mimic sharp traffic spikes (subsection 2.3).
SWE-Bench workload. We use a subset of widely-used SWE-Bench_oracle [47] targeting the sqlfluff repo [11], where the agent resolves GitHub issues. Each request represents one issue. This stresses caching because multiple issues repeatedly access shared files, such as core files and project documents.
Baseline systems. To evaluate the end-to-end performance of an agent system, we compare three primary configurations, all built atop the vLLM framework [63] for fair comparison. We compare: Agent_vanilla (no cache; every request hits the remote API), Agent_exact (traditional exact-match KV cache), Agent_Asteria (our full system), and an ablation Agent_ANN that uses only ANN similarity. As Agent_ANN is impractical for production (subsection 3.2), we restrict it to accuracy analysis (subsection 6.6).
@@ 6.2 Overall Performance On Cache Ratio
Using the setup above, we vary the cache size ratio and report throughput, cache-hit rate, and latency across three workloads.
@@
Skewed workload. We first evaluate Asteria on search benchmarks that exhibit a head–tail popularity skew, a key access pattern from subsection 2.3, to show the benefits of semantic matching. As shown in Figure 7, Asteria outperforms the baselines across throughput, cache hit rate, and latency on all four datasets (Zilliz-GPT, HotpotQA, Musique, and 2Wiki). For instance, on Musique, Agent_Asteria achieves up to a 3.6 higher throughput than Agent_exact, driven by its superior cache efficiency: Agent_Asteria sustains over 85% hit rates while the exact-match cache stays below 20%. This high hit rate reduces end-to-end latency by up to 4 and minimizes reliance on external APIs, avoiding rate limits such as the Google Search API’s 100 queries per minute cap. Note that absolute latencies exceed raw network RTTs because requests to external APIs experience queueing and backoff under rate limits, which inflates latencies, as will be shown in subsection 6.4. This performance gap stems from semantic diversity. Agent_exact fails on syntactic variations, yielding few cache hits, while Agent_Asteria groups semantically similar queries, converting semantic locality into high cache efficiency.
@@
Trend-driven workload. We next evaluate Agent_Asteria on bursty workloads synthesized from Google Trends to test its adaptability. As shown in Figure 8, it achieves up to a 3.8 throughput improvement over the Agent_vanilla baseline. This gain stems from its ability to maintain a high cache hit rate of nearly 95%, effectively absorbing the intense but temporary demand of each trending topic. This performance gain comes from the LCFU eviction policy, which is tailored for temporal dynamics. Traditional caches retain obsolete data from past trends, reducing effective capacity. In contrast, LCFU integrates an item’s staticity into its priority score. As a volatile topic’s popularity wanes, its cached entries are automatically deprioritized and become prime candidates for eviction. This self-cleaning mechanism ensures that cache space is continuously reclaimed for the next wave of content, which is crucial for maintaining high performance during bursty events.
@@
SWE-Bench workload. To demonstrate the generalizability of our approach, we evaluate Agent_Asteria on a code generation workload using the SWE-Bench. As shown in Figure 9, the results reveal a significant 20% throughput improvement over both Agent_vanilla and Agent_exact, driven by a cache hit rate approaching 45%. While more modest than the search results, these gains are highly impactful for the complex domain of software engineering and validate the versatility of our caching approach. The caching opportunity in this domain arises from shared file dependencies across tasks, such as when an agent resolves multiple GitHub issues within the same repository. A traditional cache like Agent_exact is ineffective because it treats requests for different parts of the same file as distinct misses. Agent_Asteria, however, semantically identifies the shared file context, enabling it to serve these requests from the cache. Although the inherent diversity of coding tasks leads to a lower overall hit rate compared to search, serving nearly half of all file access requests locally provides a substantial reduction in latency and system load.
@@ 6.3 Scalability Under Concurrency
To assess the Asteria’s scalability under realistic workloads, we evaluate its end-to-end throughput with varying request concurrency. We use the Musique dataset at a fixed cache ratio of 0.4, a representative setting from our prior analysis. As shown in Figure 10, the results highlight Asteria’s significant scalability advantage, delivering up to a 5.7 and 4.5 throughput improvement comparing to the baselines. While the baseline systems quickly saturate, Agent_Asteria demonstrates strong scaling, achieving a throughput of 4.89 req/s at a request rate of 8. This represents a 4.5 improvement over Agent_exact (1.09 req/s) and a 5.7 improvement over Agent_vanilla (0.86 req/s) under heavy load. The throughput for both Agent_vanilla and Agent_exact plateaus early, at just around 1 req/s. This is because their performance is fundamentally bound by the remote data retrieval. With low cache hit rates, nearly every request involves a high-latency external API call, waiting for these remote responses. Increasing concurrency merely leads to longer queues, not higher throughput. In contrast, Agent_Asteria’s throughput scales nearly linearly with concurrency up to a request rate of 8, where high cache efficiency serves most requests locally and fully utilizes GPU parallelism—driving the agent to hardware-level capacity.
@@ 6.4 Performance Breakdown
Latency analysis. To isolate pure request latency without rate-limit effects, we measure single-request breakdowns at low concurrency (Figure 11). While the core agent inference time remains constant at 0.6 s for both systems, Asteria cuts the total latency from 1.08 s in the vanilla agent down to just 0.61s. This is achieved by almost entirely eliminating the 0.48 s external retrieval bottleneck that dominates the baseline’s execution time. In its place, Asteria introduces a minimal and predictable local overhead of only 0.05 s, comprising 0.02 s for cache retrieval and 0.03 s for judger validation. This highly advantageous trade-off, which swaps a costly 0.48 s remote call for a swift 0.05 s local computation, validates our core design principle: the marginal cost of intelligent local caching is vastly outweighed by the savings from avoiding slow remote data fetches.
@@
| Agent_Asteria | 1.5 | 4.16 |
Cloud API rate limit and throughput analysis. We then study throughput under realistic load where rate limits matter (Figure 12). Beyond reducing single-request latency, Asteria significantly enhances system throughput and scalability by mitigating the critical bottleneck of external API rate limits. This is achieved by fundamentally reducing the system’s reliance on external services, as demonstrated in Figure 12. Due to high cache miss rates, the non-cached Agent_vanilla baseline generates approximately 1300 external API calls for the given task, leading to significant throttling and a high retry ratio of 25%. In stark contrast, Agent_Asteria slashes the API call count to just 103, a 92% reduction. This efficiency virtually eliminates throttling, cutting the retry ratio to a negligible 0.50%.
To quantify how this API traffic reduction translates into a throughput advantage, we run a controlled experiment as shown in Table 4. We use a self-deployed RAG data service with 300 ms here as we cannot cancel API rate limit for Google service. Even without rate limits, Asteria already provides a 1.5 throughput improvement over the vanilla system due to its inherent latency savings. However, when a realistic API rate limit is enforced, the advantage becomes far more pronounced, with Asteria achieving a 4.2 throughput gain over the throttled baseline. This comparison reveals that avoiding the rate-limiting bottleneck alone contributes an additional 2.8 improvement. By serving most requests locally, Asteria sidesteps rate-limit bottlenecks and scales robustly.
@@ 6.5 Cost Analysis
| Metric | Agent_vanilla | Asteria w/o Sharing | Asteria |
@@
Beyond performance, another critical aspect of deploying agentic systems is operational cost. We now analyze the cost-efficiency of Asteria by evaluating three configurations under peak load on the Musique dataset (similar to subsection 6.3): Agent_vanilla; Asteria w/o Sharing, which uses an extra GPU for the semantic judger; and the complete Asteria system. As detailed in Table 5, our analysis reveals that Asteria is not only faster but fundamentally more economical, delivering 6 more throughput per dollar than the vanilla baseline.
This cost advantage stems from resolving the API–compute trade-off. The Agent_vanilla system shows a total cost of $82.5 to achieve a low throughput of 0.87 req/s. The Asteria w/o Sharing configuration illustrates the pitfall of a naive caching approach. While dramatically boosting throughput to 4.79 req/s, it does so by requiring a dedicated second GPU for the semantic judger, doubling the GPU cost from $76 to $152 and making it the most expensive setup ($158.5). In contrast, our co-located design cuts API cost by 90% (to $0.64) without extra hardware, keeping total spend at $76.64 while retaining 95% of the two-GPU throughput. Thus, Asteria delivers 5.6 more performance at a similar cost, making it both fast and financially sustainable.
6.6 Accuracy Analysis
Then, we evaluate whether caching affects correctness by measuring Exact Match score [19], where the answer must match ground truth, as [32]. As shown in Figure 13, Asteria delivers efficiency without degrading accuracy. The Asteria-w/o judger baseline, which relies solely on ANN similarity, shows accuracy drops across datasets; e.g., on StrategyQA it reaches 0.69 vs. 0.79 for the baseline, showing the risk of naive semantic caching—vector similarity can return related but wrong results. In contrast, the full Agent_Asteria matches Agent_vanilla in accuracy, with the Semantic Judger validating ANN candidates and rejecting non-equivalent ones. It confirms Asteria achieves high caching efficiency without sacrificing correctness.
@@ 6.7 Deep Dive
Co-location study. Co-locating the Agent and Semantic Judger on a single A100 via CUDA MPS (80/20), at a representative cache ratio (0.6), retains 94% of dedicated throughput (2.72 vs. 2.89 req/s) with a 9.5% increase in p99 latency, indicating a near-baseline efficiency.
Recalibration overhead. We also quantify the cost of threshold recalibration for the semantic judger using HotpotQA/Musique. Compared to a variant without recalibration, throughput decreases by only 2%. This bounded cost comes from brief offline validation on small samples to keep the precision target, stabilizing accuracy under drift and preserving high-precision hits. In practice, the overhead is negligible relative to Asteria ’s gains.
@@ 7 Other Related Work
Semantic prompt caching. Recent LLm serving systems [9, 24, 65, 56] like Google Apigee [24]) match via text embeddings to bypass expensive LLM inference; ContextCache [65] caches multi-turn responses; VertorQ [56] adapts thresholds for similar prompts. However, these approaches are largely inward-facing (optimizing inference) and typically single-cloud/node, leaving the cross-region remote data bottleneck unaddressed. In contrast, Asteria targets the external-data layer, caching remote knowledge to cut latency and cost while complementing inference-side optimizations.
Traditional and multi-cloud caching. Beyond semantic schemes, caches rely on heuristics (LRU/LFU/TTL) and cache-aside patterns [62] with Redis [54] or Memcached [45]. Cross-region deployments introduce latency, consistency, and cost trade-offs. Examples include Macaron [50] (auto-tuned tiers cutting data-lake costs under write-through), EVCache [46] (eventual consistency via Kafka), Azure Cache for Redis (passive cross-region replication with client co-location), and Alluxio [38] (global namespace for on-demand hot data). These architectures assume exact-match keys and do not reason about semantic equivalence or topic dynamics. Asteria instead brings semantic-aware matching and adaptive policies to remote knowledge caching.
LLM-based agent memory. Finally, agent memory systems [67, 64, 29] improve LLM’s reasoning by retaining past context. For example, MemoryBank [67] stores conversational summaries to track entities; A-MEM [64] generalizes the memory structure to diverse tasks; Memory Sandbox [29] gives users explicit control. These systems give LLMs long-term memory to improve reasoning and choose the next action, but they don’t reduce the cost of carrying it out. Asteria instead targets the external tool call itself, making execution faster and cheaper.

### v2

@@ 5 Implementation and Discussion
Cortex is implemented atop vLLM [vllm_repo], intercepting tool calls and integrating with agent applications (e.g., Search-R1 [search_r1]). For co-location, we use CUDA MPS to let multiple processes share a single GPU. Even when local tools are low-latency, Cortex reduces cost by using semantic matching to avoid unnecessary tool calls and costs. We leave more advanced dynamic partitioning techniques (e.g., Green Contexts [cuda_green_contexts]) as future work.
Judge accuracy. Our framework bridges the gap between caching and semantic knowledge reuse by providing a unified cache abstraction for LLM agents, and it leverages powerful LLM-based techniques from prior research to ensure strong accuracy. The semantic judge can be easily fine-tuned or replaced for specific workloads, so its accuracy can be improved with minimal effort when needed. Given the rapid progress of small LLMs, we find this sufficient for practical use and view the judge as a pluggable component. In practice, Cortex maintains high precision via strong semantic models and periodic recalibration, with misses falling back to live fetches to preserve correctness.
@@ 6.1 Experimental Setup
Testbed. We emulate a realistic cross-region deployment. The LLM agent and all Cortex components run on an on-premise H100 cluster; remote data services run in a separate region. For search agent, we use the public Google Cloud Search API whose per-request average latency ranges between 300–500ms depending on response length, consistent with prior work [lin2024parrot]. For coding, we use a self-deployed FAISS [douze2024faiss]-based RAG service [douze2024faiss] with 300 ms latency.
Models and workloads. We evaluate two agentic scenarios: AI-powered search and code generation. The search agent is Search-R1-7B [jin2025search, jin2025model] (post-trained from Qwen-2.5 7B [qwen25model]); the coding agent is Qwen-3 8B [qwen3-8b]. Cortex uses Qwen-3 0.6B models for embedding and semantic judging [qwen3_embedding_06b, qwen3_reranker_06b]. Cortex’s internal embedding and semantic judging functions are powered by the lightweight 0.6B models from the Qwen-3 family [qwen3_embedding_06b, qwen3_reranker_06b].
Search workloads. Following the Google Trend access patterns in § 2.3, we construct: (i) Skewed workload: extracting the Google-like head–tail access pattern in § 2.3, we instantiate four search benchmarks–Zilliz-GPT [ziilz-gpt], HotpotQA [hotpotqa_flashrag], Musique [musique_flashrag], and 2Wiki-Multi-Hop [2wikimultihopqa]. For each dataset, we apply -means to the questions and keep 10 representative clusters, sampling 250 questions and constructing the skewed popularity per dataset (1000 total) and (ii) Trend-driven workload. We capture 12-hour Google Trends time series for four topics, map them to HotpotQA questions, and compress them into a 10-minute trace to mimic sharp traffic spikes (§ 2.3).
SWE-Bench workload. We use a subset of widely-used SWE-Bench_oracle [swebench_oracle] targeting the sqlfluff repo [sqlfluff_repo], where the agent resolves GitHub issues. Each request represents one issue. This stresses caching because multiple issues repeatedly access shared files, such as core files and project documents.
Baseline systems. To evaluate the end-to-end performance of an agent system, we compare three primary configurations, all built atop the vLLM framework [vllm_repo] for fair comparison. We compare: Agent_vanilla (no cache; every request hits the remote API), Agent_exact (traditional exact-match KV cache), Agent_Cortex (our full system), and an ablation Agent_ANN that uses only ANN similarity. As Agent_ANN is impractical for production (§ 3.2), we restrict it to accuracy analysis (§ 6.6).
@@ 6.2 Overall Performance On Cache Ratio
Using the setup above, we vary cache size ratio and report throughput, hit rate, and latency across three workloads.
@@
Skewed workload. We first evaluate Cortex on search benchmarks with head–tail popularity skew (§ 2.3) to show semantic matching benefits. As shown in Figure 7, Cortex outperforms the baselines across throughput, cache hit rate, and latency on all four datasets (Zilliz-GPT, HotpotQA, Musique, and 2Wiki). For instance, on Musique, Agent_Cortex achieves up to a 3.6 higher throughput: Agent_Cortex sustains over 85% hit rates while the exact-match cache stays below 20%. This high hit rate reduces end-to-end latency by up to 4 and minimizes reliance on external APIs, avoiding rate limits such as the Google Search API’s 100 queries per minute cap. Note that absolute latencies exceed raw network RTTs because requests to external APIs experience queueing and backoff under rate limits, which inflates latencies, as will be shown in § 6.4. This performance gap stems from semantic diversity. Agent_exact fails on syntactic variations, yielding few cache hits, while Agent_Cortex groups semantically similar queries, converting semantic locality into high cache efficiency.
@@
Trend-driven workload. We next evaluate Agent_Cortex on bursty workloads synthesized from Google Trends to test its adaptability. As shown in Figure 8, it achieves up to a 3.8 throughput improvement over the Agent_vanilla baseline. This gain stems from its ability to maintain a high cache hit rate of nearly 95%, effectively absorbing the intense but temporary demand of each trending topic. This performance gain comes from the LCFU eviction policy, which is tailored for temporal dynamics. Traditional caches retain obsolete data from past trends, reducing effective capacity. In contrast, LCFU integrates an item’s staticity into its priority score. As a volatile topic’s popularity wanes, its cached entries are automatically deprioritized and become prime candidates for eviction. This mechanism ensures that cache space is continuously reclaimed for the next wave of content, which is crucial for maintaining high performance during bursty events.
@@
SWE-Bench workload. To demonstrate the generalizability of our approach, we evaluate Agent_Cortex on a code generation workload using the SWE-Bench. As shown in Figure 9, the results reveal a significant 20% throughput improvement over both Agent_vanilla and Agent_exact, driven by a cache hit rate approaching 45%. While more modest than the search results, these gains are highly impactful for the complex domain of software engineering and validate the versatility of our caching approach. The caching opportunity in this domain arises from shared file dependencies across tasks, such as when an agent resolves multiple GitHub issues within the same repository. A traditional cache like Agent_exact is ineffective because it treats requests for different parts of the same file as distinct misses. Agent_Cortex, however, semantically identifies the shared file context, enabling it to serve these requests from the cache. Although the inherent diversity of coding tasks leads to a lower overall hit rate compared to search, serving nearly half of all file access requests locally provides a substantial reduction in latency and system load.
@@ 6.3 Scalability Under Concurrency
To assess the Cortex’s scalability under realistic workloads, we evaluate its end-to-end throughput with varying request concurrency. We use the Musique dataset at a fixed cache ratio of 0.4, a representative setting from our prior analysis. As shown in Figure 10, the results highlight Cortex’s significant scalability advantage, delivering up to a 5.7 and 4.5 throughput improvement comparing to the baselines. While the baseline systems quickly saturate, Agent_Cortex demonstrates strong scaling, achieving a throughput of 4.89 req/s at a request rate of 8. This represents a 4.5 improvement over Agent_exact (1.09 req/s) and a 5.7 improvement over Agent_vanilla (0.86 req/s) under heavy load. The throughput for both Agent_vanilla and Agent_exact plateaus early, at just around 1 req/s. This is because their performance is fundamentally bound by the remote data retrieval. With low cache hit rates, nearly every request involves a high-latency external API call, waiting for these remote responses. Increasing concurrency merely leads to longer queues, not higher throughput. In contrast, Agent_Cortex’s throughput scales nearly linearly with concurrency up to a request rate of 8, where high cache efficiency serves most requests locally and fully utilizes GPU parallelism—driving the agent to hardware-level capacity. Beyond this point, Cortex maintains stable throughput (5 req/s) even as the request rate increases to 32, with only minor degradation (3%) at the highest load due to increased memory pressure. The result demonstrates the robustness of our co-location design which provides sufficient resources for both the agent and the semantic judge without degrading to vanilla behavior under stress.
@@ 6.4 Performance Breakdown
Latency analysis. To isolate pure request latency without rate-limit effects, we measure single-request breakdowns at low concurrency (Figure 11). While agent inference remains 0.6 s, Cortex cuts the total latency from 1.08 s to just 0.61s. This is achieved by almost entirely eliminating the 0.48 s external retrieval bottleneck that dominates the baseline’s execution time. In its place, Cortex introduces a minimal and predictable local overhead of only 0.05 s, comprising 0.02 s for cache retrieval and 0.03 s for judge validation. This trade-off validates our design principle: the marginal cost of intelligent local caching (0.05 s) is vastly outweighed by the savings from avoiding slow remote data fetches (0.48 s).
@@
| Agent_Cortex | 1.5 | 4.16 |
Cloud API rate limit and throughput analysis. We then study throughput under realistic load where rate limits matter (Figure 12). Beyond reducing single-request latency, Cortex significantly enhances system throughput and scalability by mitigating the critical bottleneck of external API rate limits. This is achieved by fundamentally reducing the system’s reliance on external services, as demonstrated in Figure 12. Due to high cache miss rates, the non-cached Agent_vanilla baseline generates approximately 1300 external API calls for the given task, leading to significant throttling and a high retry ratio of 25%. In stark contrast, Agent_Cortex slashes the API call count to just 103, a 92% reduction. This efficiency virtually cuts the retry ratio to a negligible 0.50%.
To quantify how this API traffic reduction translates into a throughput advantage, we run a controlled experiment as shown in Table 4. We use a self-deployed RAG data service with 300 ms here as we cannot cancel API rate limit for Google service. Even without rate limits, Cortex already provides a 1.5 throughput improvement over the vanilla system due to its inherent latency savings. However, when a realistic API rate limit is enforced, the advantage becomes far more pronounced, with Cortex achieving a 4.2 throughput gain over the throttled baseline. This comparison reveals that avoiding the rate-limiting bottleneck alone contributes an additional 2.8 improvement. By serving most requests locally, Cortex sidesteps rate-limit bottlenecks and scales robustly.
@@ 6.5 Cost Analysis
| Metric | Agent_vanilla | Cortex w/o Sharing | Cortex |
@@
We now analyze cost-efficiency by evaluating three configurations under peak load on Musique (similar to § 6.3): Agent_vanilla; Cortex w/o Sharing, which uses an extra GPU for the semantic judge; and the complete Cortex system. As detailed in Table 5, our analysis reveals that Cortex is not only faster but fundamentally more economical, delivering 6 more throughput per dollar than the vanilla baseline.
This cost advantage stems from resolving the API–compute trade-off. The Agent_vanilla system shows a total cost of $82.5 to achieve a low throughput of 0.87 req/s. The Cortex w/o Sharing configuration illustrates the pitfall of a naive caching approach. While dramatically boosting throughput to 4.79 req/s, it does so by requiring a dedicated second GPU for the semantic judge, doubling the GPU cost from $76 to $152 and making it the most expensive setup ($158.5). In contrast, our co-located design cuts API cost by 90% (to $0.64) without extra hardware, keeping total spend at $76.64 while retaining 95% of the two-GPU throughput. Thus, Cortex delivers 5.6 more performance at a similar cost, making it both fast and financially sustainable.
@@ 6.6 Deep Dive
Accuracy Analysis We measure correctness using Exact Match score [exactmatch2025, search_r1]. As shown in Figure 13, Cortex-w/o judge shows accuracy drops across datasets (e.g., StrategyQA: 0.69 vs. 0.79), demonstrating that naive ANN similarity can return wrong results. The full Cortex matches baseline accuracy—the Semantic Judge rejects non-equivalent candidates, confirming Cortex achieves caching efficiency without sacrificing correctness.
@@
Co-location study. Co-locating the Agent and Semantic Judge on a single A100 via CUDA MPS (80/20), at a representative cache ratio (0.6), retains 94% of dedicated throughput (2.72 vs. 2.89 req/s) with a 9.5% increase in p99 latency, indicating a near-baseline efficiency.
Recalibration overhead. We also quantify the cost of threshold recalibration for the semantic judge using HotpotQA/Musique. Compared to a variant without recalibration, throughput decreases by only 2%. This bounded cost comes from brief offline validation on small samples to keep the precision target, stabilizing accuracy under drift and preserving high-precision hits. In practice, the overhead is negligible relative to Cortex ’s gains.
@@ 7 Other Related Work
Semantic prompt caching. Recent LLM serving systems [bang2023gptcache, google-cache, yan2025contextcache, schroeder2025adaptive] like Google Apigee [google-cache] match via embeddings to bypass inference; ContextCache [yan2025contextcache] caches multi-turn responses; VectorQ [schroeder2025adaptive] adapts thresholds for similar prompts. However, these approaches are largely internal to the inference process and typically for single-cloud/node, leaving the cross-region remote data bottleneck unaddressed. In contrast, Cortex operates at the external-data layer, caching knowledge retrieved via tool calls to cut cross-region latency and API costs. The two approaches are thus orthogonal and can be deployed together for complementary benefits.
Traditional and multi-cloud caching. Beyond semantic schemes, caches rely on LRU/LFU/TTL heuristics with Redis [redis-cache] or Memcached [memcached-cache]. Cross-region deployments introduce latency, consistency, and cost trade-offs. Examples include Macaron [park2024reducing] (auto-tuned tiers cutting data-lake costs under write-through), EVCache [netflix-cache] (eventual consistency via Kafka), Azure Cache for Redis (passive cross-region replication with client co-location), and Alluxio [li2018alluxio] (global namespace for on-demand hot data). These architectures assume exact-match keys, while Cortex instead brings semantic-aware matching and adaptive policies to remote knowledge caching.
LLM-based agent memory. Finally, agent memory systems [zhong2023memorybankenhancinglargelanguage, xu2025amemagenticmemoryllm, huang2023memorysandboxtransparentinteractive] improve LLM’s reasoning by retaining past context. For example, MemoryBank [zhong2023memorybankenhancinglargelanguage] stores conversational summaries to track entities; A-MEM [xu2025amemagenticmemoryllm] generalizes the memory structure to diverse tasks; Memory Sandbox [huang2023memorysandboxtransparentinteractive] gives users explicit control. These systems give LLMs long-term memory to improve reasoning and choose the next action, but they don’t reduce the cost of executing tool calls. Cortex instead targets the external tool call itself, caching external results to make agent operation faster and cheaper.